# Sampling with time-changed Markov processes

### v1

@@ 4.1. Ergodicity of time changed processes
The proof of Theorem 4.5 relies on the fact that the bounded function $\bar{\mathrm{V}}(z)=\mathrm{V}(z)/(1+\mathrm{V}(z))$ satisfies a drift condition of the form (24) for the generator of the process $Y$. This theorem can be applied to obtain uniform ergodicity of the time-changed PDMPs discussed in Section 3.3, taking advantage of the theory developed by Bierkens et al. (2019b); Deligiannidis et al. (2019); Durmus et al. (2020).
@@ 5. Connection to space transformations
(c) Space-transformed ZZP, with transformation $\mathrm{H}(y)=y(1-\lvert y\rvert^{2})^{-1/2}$ for $y\in\{z\in\mathbb{R}^{2}:\lvert z\rvert\leq 1\}.$
(d) Base process for the space-transformed ZZP, that is a standard ZZP with target $\tilde{\mu}(y,w)=\mu(\mathrm{H}(y),w)s(\mathrm{H}(y))$ with $y\in\{z\in\mathbb{R}^{2}:\lvert z\rvert\leq 1\}.$
@@ 6.1. Sampling from a mixture of normal distributions
(c) Discretisation of the time-changed process with $s(x)=1$.
@@ Acknowledgements
We would like to thank Joris Bierkens, Paul Dobson, Samuel Power, and Samuel Livingstone for helpful discussions and comments on various versions of this manuscript.
@@
The authors would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Stochastic systems for anomalous diffusion where work on this paper was undertaken. This work was supported by EPSRC (EP/Z000580/1).
@@ F.1. Eyring-Kramers formula for the time changed ZZP
We now obtain an Eyring-Kramers formula for the one dimensional, time changed ZZP in the context of Section 3.2.1. This type of result was first obtained in[Monmarché, 2016, Theorem 1.1] for the standard ZZP. We consider a target of the form $\mathrm{U}_{\varepsilon}=\mathrm{U}/\varepsilon$, where $\mathrm{U}$ is a double well potential, and we characterise the behaviour of the process as $\varepsilon\to 0$. Let $\tau:=\inf\{t>0:X_{t}=x_{1}\},$ i.e. the time elapsed before the process finally moves from its initial condition $(X_{0},V_{0})=(x_{0},-1)$ to $x_{1}$. In the next proposition we state our result for speed functions of the form $s_{\varepsilon}(x)=s(x)^{1/\varepsilon}$, where $s$ is a fixed speed function, satisfying the properties discussed in Section 3.2.1.

### v2

@@ 4.1. Ergodicity of time changed processes
The proof of Theorem 4.5 relies on the fact that the bounded function $\bar{\mathrm{V}}(z)=\mathrm{V}(z)/(1+\mathrm{V}(z))$ satisfies a drift condition of the form (22) for the generator of the process $Y$. This theorem can be applied to obtain uniform ergodicity of the time-changed PDMPs discussed in Section 3.1.2, taking advantage of the theory developed by Bierkens et al. (2019b); Deligiannidis et al. (2019); Durmus et al. (2020).
@@ 5. Connection to space transformations
(c) Space-transformed ZZP, with transformation $\mathrm{H}(y)=y(1-\lvert y\rvert^{2})^{-1/2}$ for $y\in\{z\in\mathbb{R}^{2}:\lvert z\rvert<1\}.$
(d) Base process for the space-transformed ZZP, that is a standard ZZP with target $\tilde{\mu}(y,w)=\mu(\mathrm{H}(y),w)s(\mathrm{H}(y))$ with $y\in\{z\in\mathbb{R}^{2}:\lvert z\rvert<1\}.$
@@ 6.1. Sampling from a mixture of normal distributions
(c) Discretisation without time-change ($s(x)=1$).
@@ Acknowledgements
The authors would like to thank Joris Bierkens, Paul Dobson, Pierre Monmarché, Samuel Power, and Samuel Livingstone for helpful discussions and comments on various versions of this manuscript.
@@
The authors would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Stochastic systems for anomalous diffusion where work on this paper was undertaken. This work was supported by EPSRC (EP/Z000580/1). The authors would also like to thank Codina Cotar for inviting them to participate at the INI programme.
@@ F.1. Eyring-Kramers formula for the time changed ZZP
We now obtain an Eyring-Kramers formula for the one dimensional, time changed ZZP in the context of Section 3.1. This type of result was first obtained in[Monmarché, 2016, Theorem 1.1] for the standard ZZP. We consider a target of the form $\mathrm{U}_{\varepsilon}=\mathrm{U}/\varepsilon$, where $\mathrm{U}$ is a double well potential, and we characterise the behaviour of the process as $\varepsilon\to 0$. Let $\tau:=\inf\{t>0:X_{t}=x_{1}\},$ i.e. the time elapsed before the process finally moves from its initial condition $(X_{0},V_{0})=(x_{0},-1)$ to $x_{1}$. In the next proposition we state our result for speed functions of the form $s_{\varepsilon}(x)=s(x)^{1/\varepsilon}$, where $s$ is a fixed speed function, satisfying the properties discussed in Section 3.1.